# FOOL: Addressing the Downlink Bottleneck in Satellite Computing with Neural Feature Compression

### v1

@@ 7.2 Reliance on Foundational Models
Earth Observation (EO) requires considerations not included in common object detection objectives and architectural components, and widespread foundational models for satellite imagery have yet to emerge. Nonetheless, proprietary offerings already exist [83], and we argue that the community drives to open-source solutions will inevitably mitigate the limitation. We worked around not having access to an EO-native foundational model by only freezing the shallow layers to train the predictors. In other words, each predictor is complemented with a large backbone trained for a certain GSD range, i.e., not preserving the details with adequate granularity would result in predictive loss. Consequently, applying the saliency-guided training strategy of FrankenSplit [17] for tasks with varying GSD was difficult, as it relies on “soft signals” from deeper layers to regularize the distortion measure. For example, we can extract saliency maps to train an SVBI codec that reliably results in lossless prediction for coarser-grained terrestrial and aerial imagery, such as COCO [84], DOTA-2 [64], and RarePlanes [72]. However, it would struggle with the finer-grained xView task. As a side note, this was the initial motivation of the keypoint modality, i.e., a replacement task-agnostic method to lower the bitrate without predictive loss. The advantage of using keypoints is that it completely removes any dependency on deeper layers. In contrast, the saliency-guided strategy can reduce bitrate without any runtime overhead, as it is a training objective regularizer.
@@ 8 Conclusions
This work introduced a novel compression method that addresses the downlink bottleneck in LEO without relying on prior knowledge of downstream tasks. A rigorous evaluation showed that FOOL increases data volume with advancements that, to the best of our knowledge, are unprecedented. The rate gains are primarily from the task-agnostic context. Additionally, the AIT layers further improve compression performance with an overhead that does not outweigh the processing throughput gains from batch parallelization. Lastly, we transparently listed limitations that future work should consider and identified promising future research directions for OEC based on novel insights.

### v2

@@ 7.2 Reliance on Foundational Models
Earth Observation (EO) requires considerations not included in common object detection objectives and architectural components, and widespread foundational models for satellite imagery have yet to emerge. Nonetheless, proprietary offerings already exist [81], and we argue that the community drives to open-source solutions will inevitably mitigate the limitation. We worked around not having access to an EO-native foundational model by only freezing the shallow layers to train the predictors. In other words, each predictor is complemented with a large backbone trained for a certain GSD range, i.e., not preserving the details with adequate granularity would result in predictive loss.
@@ 8 Conclusions
This work introduced a novel compression method that addresses the downlink bottleneck in LEO without relying on prior knowledge of downstream tasks. A rigorous evaluation showed that FOOL increases data volume with advancements that, to the best of our knowledge, are unprecedented. The rate reductions are primarily from the task-agnostic context. Additionally, the ITA layers further improve compression performance with an overhead that does not outweigh the processing throughput gains from batch parallelization. Lastly, we transparently listed limitations that future work should consider and identified promising future research directions for OEC based on novel insights.